OpenStudy (anonymous):

Simplify -3(p-5pp) + 3pp-6P

OpenStudy (anonymous):

-3(p-5pp) + 3pp-6P = -3p+15pp+3pp-6p =-9p+18pp

OpenStudy (anonymous):

if u wanna continue = -9(p-2pp)